### v1

@@ 7 Experiments
As discussed in Section 1, we attempt to optimize client-side TFHE operations and communication costs. This section answers the following questions via extensive empirical evaluations to confirm how much these requirements are satisfied.
@@ 1. Q1:
    How fast is the running time on the SBC and the server?
@@ 2. Q2:
    How large is the ciphertext size exchanged between the SBC and the server?
@@ 3. Q3:
    How does the power consumption of TFHE-SBC in client-side operation compare to other methods?
The parameters for TLWE and TRLWE are the values in Table 3, which are the same as lvl1param in the TFHEpp library. This value achieves 128-bit security. All of our experiments are conducted on Raspberry Pi Zero 2W with an ARM 1GHz 4-core CPU and 512MB RAM as SBC and a machine with an Intel 1.9GHz 56-core CPU and 512GB RAM as the cloud server. We implemented TFHE-SBC using C++17 and complied with GCC-10.2.1. Our implementation is available on https://anonymous.4open.science/r/TFHE-SBC-A178.
For Q1 and Q2, we compare the encryption time on the SBC and ciphertext size of TLWE and TRLWE encryption and show the conversion time from TRLWE ciphertext to TLWE ciphertext on the server. We also compare the running time of TRLWE encryption for each condition of random number generation and PolyMul methods and discuss the optimal combination to be employed in TFHE-SBC. We also report memory usage for each configuration.
For Q3, we connect the UM24 USB meter between the power supply and the device to measure the power consumption.
@@ 7.1 Experimental Results
To answer Q1, we first evaluate the computational cost on the SBC side and then on the server side. We also show the optimization effects in TRLWE encryption in detail for the uniform and Gaussian noise generation and PolyMul, respectively.
End-to-end computation costs on SBCs. Table 1 compares the execution time, ciphertext size, and energy consumption of TLWE encryption (tlweSymEncrypt function) and TRLWE encryption (tlweSymEncrypt function) by the TFHEpp library as the baseline with TFHE-SBC. The plaintext size is set from 256 bits to 8192 bits. TFHE-SBC library runtime is the value when implemented with the fastest combination of random number generation and polynomial multiplication as described in Section 6.1. TFHEpp in environments that do not support AES accelerators is implemented using /dev/urandom as CSPRNG and the polar method for Gaussian noise. In terms of execution time, TRLWE encryption encrypts $N$ bits at a time, so the encryption times for 256-bit and 1024-bit encryption are the same. The optimized TRLWE with TFHE-SBC is approximately 15 times faster than TRLWE encryption with TFHEpp library and approximately 2490 times faster than TLWE encryption. For RAM consumption, TFHE-SBC is 12.8 times more efficient for 1024-bit encryption than baseline TLWE encryption. For TRLWE, memory usage was 336.2 KiB for our method, 1.01 times the baseline. The increase of 4 KiB memory usage in TFHE-SBC is due to the random number generation being replaced by BLAKE2, as can be observed from Table 6. This means that TFHE-SBC achieved 15 times faster encryption without much increase in memory usage from non-optimized TRLWE.
@@
Figure 3: Performance of TRLWEtoTLWEs. It is clear that end-to-end encryption of TFHE-SBC is faster than TLWE encryption on the client, even when conversion time is taken into account.
@@
Switching from TRLWE to TLWE. We evaluate the server-side computational cost of the TFHE-SBC adapter library. Note that if the client is TLWE encrypted, such costs are ignored. Figure 3(a) shows the execution time of TRLWEtoTLWEs on the server side with TFHE-SBC’s adapter library. The execution is single-threaded. For a 1024-bit plaintext, TRLWE encryption with the TFHE-SBC’s device library requires 3.1 ms on the client device and 1.0 ms for conversion on the server, for a total of 4.1 ms. In this case, TLWE encryption requires 7705.6 ms from Table 1. Therefore, end-to-end encryption of TFHE-SBC is faster than TLWE encryption on the client, even when the total execution time of TRLWEtoTLWEs is taken into account. Figure 3(b) shows that the peak memory usage of TRLWEtoTLWE increases linearly. TRLWEtoTLWE requires 8KiB TRLWE ciphertext as input and outputs 4MiB TLWE ciphertexts, so at least 4plaintext $4*\frac{\text{plaintext size}}{N}$ MiB is reserved. Additional memory is required to concatenate the output TLWE by std::copy.
@@
Comparison of the noise sampling. From Table 6, changing from /dev/urandom to BLAKE2 as the CSPRNG had the greatest speedup effect. Since CSPRNG is involved in generating random numbers sampled from uniform and Gaussian distributions, it indicates that an improvement in the speed of pseudorandom number generation affects the speed of sampling from each distribution. BLAKE2 generates random numbers 8 times faster than /dev/urandom for both uniform and Gaussian distributions with an additional 4 KiB increase in memory usage. As shown in Section 6.1 and Table 6, the execution time of Gaussian number generation improved by employing the Ziggurat method rather than the polar method employed in TFHEpp. When implemented in the same CSPRNG, Gaussian sampling using the Ziggurat method is more than twice as fast as Gaussian sampling using the polar method. These results are derived from the difference in rejection probabilities of each method.
@@
Comparison of the polynomial multiplication. We test the performance of PolyMul with two configurations, as shown in Table 7. In the “on-the-fly” version, we execute FFTs for the secret key and the uniform noise, so the FFT is performed twice for each encryption. In the “reuse” flavor, the FFT value of the secret key is precomputed. It requires less computation than the “on-the-fly” case since the FFT for the secret key only needs to be calculated once per encryption. Peak memory usage is the same for both cases. PolyMul “re-use” case, which eliminates one FFT execution, improves the computational cost of the PolyMul “on-the-fly” case by 28%.
@@
To answer Q2, Table 1 shows the communication costs between SBCs and the server. The communication cost here refers to the ciphertext size. The ciphertext size of TLWE, which encrypts one bit at a time, is approximately 512 times larger than that of TRLWE, which encrypts $N$ bits at once. This result is consistent with the theoretical result in Section 4.2. That is, generating TRLWE ciphertext leads to a reduction in the amount of client-server communication.
@@
To answer Q3, we conduct measurements of the SBC’s power consumption. The rightmost column of Table 1 shows the energy consumed in the end-to-end encryption for each plaintext size on SBC. The idle power consumption in Raspberry Pi Zero 2W is 600 mW. TFHE-SBC consumes 5 mJ per encryption, nearly 2000 times less energy than TLWE encryption. In addition, our approach achieves 12 times more energy efficiency than baseline TRLWE encryption. This is because TFHE-SBC has a better encryption runtime.
8 Limitations and Conclusions
We proposed TFHE-SBC, the first TFHE scheme implementation of the client-side operations for SBCs. TFHE-SBC archive the efficient TLWE encryption and reduce the ciphertext size through the TRLWE encryption on SBCs and its conversion to TLWE ciphertext by the server. Furthermore, we built efficient TRLWE encryption with the limited computational resources of SBC and investigated optimal random number generation and polynomial multiplication, including the memory reuse strategy. We also discussed the applicability of GPUs on Raspberry Pi in the TFHE scheme and showed the lack of accuracy of polynomial multiplication. Experimental results show that TFHE-SBC achieved 15 to 2486 times faster encryption, 512 times more efficient communication, up to 12.8 times more memory efficiency, and 12 to 2004 times more energy efficiency than the state-of-the-art on Raspberry Pi Zero 2W. Note that our framework only supports private key encryption, and public key encryption is not implemented yet. It is an interesting challenge to efficiently implement encryption on SBCs that requires a larger number of noise samplings and a much larger public key than the secret key.

### v2

@@ 7 Experiments
As established in Section 1, our primary objectives are to optimize client-side TFHE operations and reduce communication costs. This section addresses the following research questions through comprehensive empirical evaluations to quantify the extent to which our solution meets these requirements:
@@ 1. Q1:
    How does TFHE-SBC improve computational efficiency on both the SBC client and server?
@@ 2. Q2:
    How does TFHE-SBC reduce the communication overhead between the SBC and the server?
@@ 3. Q3:
    How does the energy efficiency of TFHE-SBC’s client-side operations compare to alternative methods?
For our experiments, we use the parameter values for TLWE and TRLWE as specified in Table 3, which correspond to the lvl1param settings in the TFHEpp library. These parameters achieve 128-bit security. All experiments are conducted on a Raspberry Pi Zero 2W (ARM 1GHz 4-core CPU, 512MB RAM) as the SBC client and a server machine (Intel 1.9GHz 56-core CPU, 512GB RAM) as the cloud server. We implemented TFHE-SBC in C++17 and compiled it with GCC-10.2.1. Our implementation is available at https://anonymous.4open.science/r/TFHE-SBC-A178.
To address Q1 and Q2, we evaluate and compare the encryption time on the SBC and ciphertext size for both TLWE and TRLWE encryption methods. We also measure the conversion time from TRLWE ciphertext to TLWE ciphertext on the server. Additionally, we analyze the execution time of TRLWE encryption across different random number generation and polynomial multiplication methods to determine the optimal configuration for TFHE-SBC. We report memory utilization metrics for each configuration to provide a comprehensive performance assessment.
For Q3, we employ a UM24 USB power meter connected between the power supply and the device to precisely measure energy consumption during client-side operations.
@@ 7.1 Experimental Results
To address Q1, we first evaluate the computational efficiency on the SBC side and then on the server side. We further analyze specific optimization effects in TRLWE encryption for uniform and Gaussian noise generation as well as polynomial multiplication.
End-to-end computational efficiency on SBCs. Table 1 presents a comparison of execution time, ciphertext size, and energy consumption across TLWE encryption (tlweSymEncrypt function) and TRLWE encryption (trlweSymEncrypt function) by the TFHEpp library as baselines, and TFHE-SBC. TFHE-SBC measurements represent the optimal combination of random number generation and polynomial multiplication as described in Section 6.2. TFHEpp in environments without AES accelerators uses /dev/urandom as CSPRNG and the polar method for Gaussian noise. Regarding execution time, since TRLWE encryption processes $N$ bits at a time, the encryption times for 256-bit and 1024-bit plaintexts are the same. The optimized TRLWE with TFHE-SBC is approximately 15 times faster than TRLWE encryption with TFHEpp library and approximately 2490 times faster than TLWE encryption. For RAM consumption, TFHE-SBC is 12.8 times more efficient for 1024-bit encryption than baseline TLWE encryption. For TRLWE, memory usage was 336.2 KiB for our method, 1.01 times the baseline. The 4 KiB increase in memory usage in TFHE-SBC is due to the random number generation being replaced by BLAKE2, as evidenced in Table 6. This demonstrates that TFHE-SBC achieved 15 times faster encryption without much increase in memory usage from non-optimized TRLWE.
@@
Figure 3: Performance of TRLWEtoTLWEs. Results confirm TFHE-SBC’s end-to-end encryption outperforms direct TLWE encryption, even with conversion overhead.
@@
TRLWE to TLWE conversion performance. We evaluate the server-side computational overhead of the TFHE-SBC adapter library. Note that if the client is TLWE encrypted, such costs are ignored. Figure 3(a) shows the execution time of TRLWEtoTLWEs on the server side with TFHE-SBC’s adapter library. The execution is single-threaded. For a 1024-bit plaintext, TRLWE encryption with the TFHE-SBC’s device library requires 3.1 ms on the client device and 1.0 ms for conversion on the server, for a total of 4.1 ms. In this case, TLWE encryption requires 7705.6 ms from Table 1. Therefore, end-to-end encryption of TFHE-SBC is faster than TLWE encryption on the client, even when the total execution time of TRLWEtoTLWEs is taken into account. Figure 3(b) shows that the peak memory usage of TRLWEtoTLWE increases linearly. TRLWEtoTLWE requires 8KiB TRLWE ciphertext as input and outputs 4MiB TLWE ciphertexts, so at least 4plaintext $4*\frac{\text{plaintext size}}{N}$ MiB is reserved. Additional memory is required to concatenate the output TLWE by std::copy.
@@
Noise sampling optimization analysis. From Table 6, changing from /dev/urandom to BLAKE2 as the CSPRNG had the greatest speedup effect. Since CSPRNG is involved in generating random numbers sampled from uniform and Gaussian distributions, enhancing pseudorandom number generation directly improves sampling performance. BLAKE2 generates random numbers 8 times faster than /dev/urandom for both distributions with an additional 4 KiB increase in memory usage. As detailed in Section 6.2 and Table 6, the Ziggurat method delivers substantial improvements over the polar method employed in TFHEpp. When implemented with the same CSPRNG, Gaussian sampling using the Ziggurat method is more than twice as fast as the polar method. This performance differential stems from the difference in rejection probabilities of each method.
@@
Polynomial multiplication optimization. We test the performance of PolyMul with two configurations, as shown in Table 7. In the “on-the-fly” version, we execute FFTs for the secret key and the uniform noise, so the FFT is performed twice for each encryption. In the “reuse” flavor, the FFT value of the secret key is precomputed. It requires less computation than the “on-the-fly” case since the FFT for the secret key only needs to be calculated once per encryption. Peak memory usage is the same for both cases. PolyMul “reuse” case, which eliminates one FFT execution, improves the computational cost of the PolyMul “on-the-fly” case by 28%.
@@
To address Q2, Table 1 shows the communication costs between SBCs and the server. The communication cost here refers to the ciphertext size. The ciphertext size of TLWE, which encrypts one bit at a time, is approximately 512 times larger than that of TRLWE, which encrypts $N$ bits at once. This result is consistent with the theoretical result in Section 5.1. These findings confirm that generating TRLWE ciphertext leads to a reduction in the amount of client-server communication.
@@
To address Q3, we conduct measurements of the SBC’s power consumption. The rightmost column of Table 1 shows the energy consumed in the end-to-end encryption for each plaintext size on SBC. The idle power consumption in Raspberry Pi Zero 2W is 600 mW. TFHE-SBC consumes 5 mJ per encryption, nearly 2000 times less energy than TLWE encryption. In addition, our approach achieves 12 times more energy efficiency than baseline TRLWE encryption. This substantial improvement directly correlates with TFHE-SBC’s enhanced encryption runtime.
8 Conclusion and Future Work
We presented TFHE-SBC, the first implementation of TFHE client-side operations designed specifically for resource-constrained SBCs. Our framework achieves significant improvements in both computational efficiency and communication overhead by employing TRLWE encryption on SBCs with server-side conversion to TLWE ciphertext. We developed key optimizations tailored to SBCs’ limited resources, including enhanced random number generation and efficient polynomial multiplication with memory reuse. Our investigation of GPU acceleration on the Raspberry Pi identified critical accuracy limitations that informed our CPU-focused design.
Experimental evaluations demonstrate that TFHE-SBC outperforms state-of-the-art implementations on the Raspberry Pi Zero 2W across all metrics: achieving encryption speeds 15 to 2486 times faster, communication efficiency 512 times greater, memory utilization up to 12.8 times more efficient, and energy consumption 12 to 2004 times lower than existing approaches.
While TFHE-SBC makes TFHE more practical for IoT applications, our current implementation supports only private key encryption. Extending the framework to support public key encryption presents additional challenges, as it requires substantially more noise sampling operations and memory resources. Addressing these challenges remains an important direction for future research.